# Attacking the Spike: On the Transferability and Security of Spiking Neural Networks to Adversarial Examples

### v3

@@ Appendix D Auto-SAGA Supplementary Material
    For single MIM and PGD attacks, we use $\epsilon_{step}=0.01$, attack step = 40 to generate AEs from each model, and list the highest attack success rate on each pair of the models.
@@
    For basic SAGA, we set the attack as a balanced version of SAGA that uses coefficients $\alpha_{1}=\alpha_{2}=0.5$ for two models to generate AEs and get the attack success rate among both models.
  3. 3.
    As for Auto-SAGA, we set learning rate $r=10,000$ for the coefficients. We set attack step = 40 and $\epsilon_{step}=0.005$ to generate AEs.
Model 1 Model 2 Max MIM Max PGD Max Auto Basic SAGA Auto-SAGA
C-V S-R-BP 40.7% 33.6% 40.4% 49.8% 93.2%
C-V S-V-BP 59.4% 51.3% 57.6% 67.2% 94.7%
C-V S-V-T10 73.1% 68.6% 70.3% 78.6% 84.0%
C-V S-R-T10 69.6% 46.6% 68.5% 84.4% 91.8%
S-R-BP S-V-T10 41.7% 33.7% 29.8% 45.3% 64.3%
V-L16 S-R-BP 28.3% 23.5% 22.1% 74.5% 78.9%
V-L16 S-V-BP 33.9% 20.3% 18.8% 70.0% 85.4%
V-L16 S-V-T10 25.7% 15.3% 13.6% 33.0% 91.5%
V-L16 S-R-T10 27.2% 17.4% 15.3% 60.8% 93.8%
C-101x3 S-R-BP 38.3% 32.6% 30.3% 52.0% 77.3%
C-101x3 S-V-BP 22.7% 16.9% 16.1% 57.0% 83.8%
C-101x3 S-V-T10 24.6% 20.3% 17.9% 44.5% 84.5%
C-101x3 S-R-T10 25.2% 21.0% 19.5% 85.8% 97.0%
Table 16: Max MIM, PGD, and Auto represent the max suc- cess rate using adversarial examples generated by model 1 and model 2 for CIFAR-100 task
Model 1 Model 2 Max MIM Max PGD Max Auto Basic SAGA Auto-SAGA
TIC-R19 HIRE-V11 68.5% 69.0% 66.1% 15.2% 72.9%
FAT-R18 HIRE-V11 28.3% 29.6% 28.5% 21.7% 50.1%
FAT-R18 TIC-R19 25.2% 25.5% 27.1% 23.8% 46.6%
DM-R18 HIRE-V11 14.8% 16.0% 17.7% 14.6% 38.0%
DM-R18 FAT-R18 27.9% 27.7% 29.5% 29.7% 37.4%
DM-R18 TIC-R19 12.1% 11.2% 11.5% 11.7% 35.8%
Table 17: Max MIM, PGD, and Auto represent the max success rate using adversarial examples generated by adversarial trained SNN model 1 and model 2 for CIFAR-100 task.
Appendix E SNN Transferability Study Supplementary Material

### v4

@@ 4 SNN Transferability Study
    How vulnerable are SNNs to adversarial examples generated from other machine learning models like Vision Transformers and CNNs?
@@
    Do non-SNN models misclassify adversarial examples created from different types of SNNs?
Formally, transferability is the phenomenon that occurs when adversarial examples generated using one model are also misclassified by a different model. Transferability studies have been done with CNNs [28, 54] and with ViTs [24]. To the best of our knowledge, the analysis of the transferability of adversarial examples with respect to SNNs has never been done. Both transfer questions posed at the start of this section, are important from a security perspective. If adversarial samples do not transfer in either direction, then either new SNN/CNN/ViT ensemble defenses are possible. In addition, if adversarial samples between different model types exhibit low transferability, new white-box attacks must be developed to be able to successfully attack both SNNs and non-SNNs simultaneously.
We briefly define how the transferability between different models is measured. Consider a white-box attack $A$ on classifier $C_{i}$ which produces adversarial example: $x_{adv}=A_{C_{i}}(x,t)$}, where $x$ is the original clean example and $t$ is the corresponding correct class label. Now consider a second classifier $C_{j}$ independent from classifier $C_{i}$. The adversarial example $x_{adv}$ transfers from $C_{i}$ to $C_{j}$ if and only if the original clean example $x$ is correctly identified by $C_{j}$ and $x_{adv}$ is misclassified by $C_{j}$:
$\{C_{j}(x)=t\}\land\{C_{j}(x_{adv})\neq t\}$ (24)
From Equation 25, we can see that a high transferability suggests models share a security vulnerability, that is, most of the adversarial examples generated by $A_{C_{i}}$ are misclassified by both models $C_{i}$ and $C_{j}$.
4.1 Transferability Experiment and Analyses
Experimental Setup: For our transferability experiment, we analyze four common white-box adversarial attacks which have been experimentally verified to exhibit transferability [33, 55]. The four attacks are the Fast Gradient Sign Method (FGSM) [18], Projected Gradient Descent (PGD) [23] the Momentum Iterative Method (MIM) [22] and Auto-PGD [25]. For each attack, we use the $l_{\infty}$ norm with $\epsilon=0.031$. For brevity, we only list the main attack parameters here and give detailed descriptions of the attacks in the appendix. When running the attacks on SNN models, we use the surrogate gradient function that worked most effectively across all the SNN models for the CIFAR datasets (Arctan) as demonstrated in Section 3. In terms of datasets, we show results for CIFAR-10. When running the transferability experiment between two models, we randomly select 1000 clean examples that are correctly identified by both models and class-wise balanced.
Models: To study the transferability of SNNs in relation to other models, we use a wide range of classifiers. These include Vision Transformers: ViT-B-32, ViT-B-16 and ViT-L-16 [56]. We also employ a diverse group of CNNs: VGG-16 [57], ResNet-20 [58] and BiT-101x3 [59]. For SNNs, we use both BP and Transfer trained models. For BP SNNs, we experiment with BP SNN VGG-16 [9] and SEW ResNet [10]. For Transfer based SNNs we study an SNN VGG-16 [7]. All the SNNs are trained following the conventional settings based on their papers. The timesteps we use for different models and corresponding clean accuracies for CIFAR-10 are given in Table 10.